# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

### 05 Aug 2021 v1

@@ 2.2 Hardness in three dimensions
The reverse direction is a little more involved. Let $\varphi$ denote a geometric embedding of $C$. Note that the tunnel of each pseudoline forms a closed topological sphere; all of which are pairwise disjoint. By a generalization of the Jordan curve theorem (also known as the Jordan–Brouwer separation theorem) any topological embedding of a (d1)(d-1)-sphere in $\mathbb{R}^{d}$ splits the space into two components; we refer to the bounded component as its inside and to the unbounded as its outside. Note that the apex $u$ can lie inside at most one tunnel; otherwise, among all tunnels containing $u$, the innermost separates $u$ from the outermost. Consequently, for at least one one copy of $\overline{C}$ contained in $C$, the apex $u$ lies in the outside of each tunnel in $\overline{C}$. From now on, we focus on the geometric embedding of this $\overline{C}$ and do not make further use of the other copy.
As a crucial property, we argue that the special edge of each pseudoline lies in the interior of its tunnel. To this end, consider $s_{i}$ and the tunnel forming a closed topological sphere incident to $s_{i}$. Note that the triangles formed by the apex $u$ and objects of the last section of the tunnel (non-incident to $s_{i}$) form a tent on top of the tunnel-sphere, see Figure 3(c). In particular, the tent together with the tunnel triangles incident to $s_{i}$ form another sphere incident to $s_{i}$. If $e_{i}$ would start towards the outside of the tunnel, then it would be trapped inside the tent-sphere which is not incident to $t_{i}$. Consequently, $e_{i}$ lies inside the tunnel. Moreover, the edges $\{u,s_{i}\}$ and $\{u,t_{i}\}$ ensure that $e_{i}$ essentially stretches over the entire tunnel. Most importantly, $e_{i}$ runs through every section of the tunnel that belongs to a crossing.
@@
Now, we consider a small ball around the apex which does not contain any other vertex, see Figure 5(a). For each triangle containing the apex $u$, the intersection with the sphere $S$ yields an arc of a great circle. Consequently, we obtain a crossing-free drawing of the neighborhood complex $\overline{C}_{u}$ of $u$ with arcs of great circles on $S$. Because $\overline{C}_{u}$ is 3-connected, there exists a unique planar drawing on $S$. Recall that $\overline{C}_{u}$ captures the crossings of the tunnels and thus the combinatorics of $L^{*}$. Moreover, for each special edge $e_{i}$, we consider the projection $a_{i}$ of the (artificial) triangle $\{s_{i},t_{i},u\}$ onto the sphere $S$; note that $\{s_{i},t_{i},u\}$ is not a triangle of $C$, see Figure 5(b). The resulting arc $a_{i}$ lies within projection of the tunnel because $e_{i}$ is contained in the tunnel. Hence, the arc arrangement has the same combinatorics/intersection pattern as $L$, i.e., the order of intersections along each arc/pseudoline are the same.
@@
Finally, we consider the supporting planes of the arcs (through $u$). Specifically, we consider the plane $p$ through the arc $a_{0}$ and apex $u$ and assume without loss of generality that $p$ is horizontal. Clearly, all arcs start to the same hemisphere of $S$ cut by $p$; let suppose the bottom hemisphere. Then, the intersection of any two arcs $a_{i}$ and $a_{j}$ lie on the bottom hemisphere; otherwise the arcs contain both intersections of the supporting great circles. Let $p^{\prime}$ denote the plane obtained by shifting $p$ slightly to the bottom. Consequently, we can ’project’ the arc arrangement onto the plane $p^{\prime}$ by considering the intersection of $p^{\prime}$ with the plane supporting the arc $a_{i}$. Clearly, this intersection yields a line. Hence, we obtain a line arrangement in the plane $p^{\prime}$ that proves the stretchability of $L$. ∎
@@
The constructed 2-complex in the proof of Lemma 4 was not pure. Specifically, the special edges are not contained in any triangle. We can obtain a pure 2-complex by adding one new vertex to each special edge such that it forms a special triangle. Similarly, we can add a private vertex to each triangle to form a pure 3-complex $C^{\prime}$. Given a geometric embedding of $C$, these new vertices can easily be added close enough to their defining set in $C$. Hence, $C$ has a geometric embedding if and only if $C^{\prime}$ has a geometric embedding in $\mathbb{R}^{3}$. Together with Lemma 4, this very small modification shows hardness of $\textsc{GEM}_{3\rightarrow 3}$.

### 05 Nov 2021 v2

@@ 2.2 Hardness in three dimensions
By definition, each tunnel forms a closed topological sphere; all of which are pairwise disjoint. By a generalization of the Jordan curve theorem, also known as the Jordan–Brouwer separation theorem, any topological embedding of a (d1)(d-1)-sphere in $\mathbb{R}^{d}$ splits the space into two components [undefaal]; we refer to the bounded component of a tunnel or any other topological sphere as its inside and to the unbounded component as its outside.
Claim 4.
There exists a copy $\overline{C}$ in $C$ such that the apex $u$ lies outside all tunnels of $\varphi(\overline{C})$.
Proof.
Note that the apex $u$ lies inside at most one tunnel of $C$; otherwise, among all tunnels containing $u$, the innermost separates $u$ from the outermost. A contradiction to the fact that shares and edge with the vertices of all tunnels. Consequently, for at least one copy of $\overline{C}$ contained in $C$, the apex $u$ lies outside of each tunnel. ∎
From now on, we focus on the geometric embedding $\varphi(\overline{C})$ of this $\overline{C}$ and do not make further use of the other copy.
Claim 5.
In $\varphi(\overline{C})$, the special edge $e_{i}$ lies inside its tunnel $T^{+}_{i}$ for all $i$.
Proof.
Consider the vertex $s_{i}$ of $e_{i}$ and the tunnel $T^{+}_{i}$ forming a closed topological sphere incident to $s_{i}$. By construction, all roof edges of the first section of $T^{+}_{i}$ are visible by the apex $u$ and thus form triangles with $u$. Note that the edges non-incident to $s_{i}$ contain a cycle, see Figure 9(a).
(a) A top view of first section.
(b) The tent over $s_{i}$.
Figure 9: Illustration for the proof of Claim 5: the spheres surrounding $s_{i}$.
The triangles between the apex $u$ and this cycle form a tent on top of $T^{+}_{i}$, see Figure 9(b). In particular, the tent together with the tunnel roof of the first section form another sphere incident to $s_{i}$. By Claim 4, $u$ is outside the tunnel. If $e_{i}$ started towards the outside of the tunnel, then it would be trapped inside the tent-sphere which is not incident to $t_{i}$. Consequently, $e_{i}$ lies inside the tunnel. ∎
Consider a special edge $e_{i}$ and the wedge $W_{i}$ defined by rays originating at $u$ through points in $e_{i}$, see Figure 10. We say a point $p$ of $(W_{i}\setminus e_{i})$ lies above $e_{i}$ if the segment $up$ does not intersect $e_{i}$; we write $p>e_{i}$. Similarly, $p$ lies below $e$ if the segment $up$ does intersect $e_{i}$; we write $p<e_{i}$.
Figure 10: The wedge $W_{i}$ of $e_{i}$ containing a point $p$ above $e_{i}$ and a point $q$ below $e_{i}$.
Claim 6.
Consider a special edge $e_{i}$. For every two points $p,q$ on $W_{i}\setminus e_{i}$ that belong to a common triangle in $T^{+}_{i}$, the above/below-relation is consistent, i.e., $p<e_{i}\iff q<e_{i}$.
Proof.
Suppose for a contradiction that $p>e_{i}$ and $q<e_{i}$. By visibility of $s_{i}$ and $t_{i}$, $p$ lies neither on the segment $us_{i}$ nor on the segment $ut_{i}$. Hence, the segment $pq$ intersects $e_{i}$ in an inner point, i.e., it intersects $e-\{s_{i},t_{i}\}$, see Figure 10. However, no triangle of $C$ contains $e_{i}$ and hence $e_{i}-\{s_{i},t_{i}\}$ and $pq$ must have an empty intersection. A contradiction. ∎
Now, we consider a small sphere $S$ around the apex $u$ which has no other vertex inside, see Figure 11(a). For each triangle in $\varphi(\overline{C})$ containing $u$, the intersection with the sphere $S$ yields an arc of a great circle. Consequently, we obtain a crossing-free drawing $D^{+}$ of the essentially 3-connected planar graph $G^{+}$ (the neighborhood complex of $u$) with arcs of great circles on $S$. Moreover, for each special edge $e_{i}$, we consider the projection $a_{i}$ of the (artificial) triangle $\{s_{i},t_{i},u\}$ onto the sphere $S$; note that $\{s_{i},t_{i},u\}$ is not a triangle of the complex, see Figure 11(b).
@@
This yields a set of arcs $\mathcal{A}$. In the remainder, we show that $\mathcal{A}$ has the same combinatorics/intersection pattern as the pseudoline arrangement $L^{*}$, i.e., the order of intersections along each arc/pseudoline are the same. Afterwards, we project $\mathcal{A}$ to a plane as illustrated in Figure 11(c) and obtain a line arrangement proving that $L^{*}$ (and thus $L$) is stretchable.
Let $D$ denote the restriction of the drawing $D^{+}$ to the graph $G$. By Claim 2, all crossing-free drawings of $G^{+}$ on the sphere are equivalent. This implies the following fact:
Claim 7.
In $D^{+}$, the helper triangle separates the subdrawing $D$ from the vertices $\{s_{j},t_{j}\mid j=0,\dots,n\}$. In particular, by convexity, the convex hull of any vertex subset $U$ in $D$ is contained in the helper triangle.
For a tunnel, we call the intersection of two consecutive sections a tunnel loop. Note that each tunnel loop consists of three edges, two of which belong to the roof of the tunnel and form a rafter; the remaining edge belongs to the bottom of the tunnel.
Claim 8.
$D\cup\mathcal{A}$ fulfills the following properties for all $i$:
  1. i)
    The arc $a_{i}$ intersects each rafter of $T_{i}$ exactly one time.
  2. ii)
    The intersection points appear in the correct order along $a_{i}$, i.e., they follow the natural order along the tunnel.
  3. iii)
    The faces of a roof section in $T_{i}$ cover an interval of $a_{i}$.
Proof.
Consider a tunnel loop o{o} of tunnel $T^{+}_{i}$ and its (projected) vertices in $D$. We first show that the arc $a_{i}$ intersects at least one (potentially artificial) segment connecting the vertices of o{o} in $D$. To this end, we may assume that $o$ is disjoint from $s_{i}$ and $t_{i}$ and consider the restriction of $\varphi(\overline{C})$ to the plane $P_{i}$ containing $u$ and $e_{i}$. By Claims 4 and 5, the apex $u$ is outside of the tunnel $T^{+}_{i}$ while $e_{i}$ is inside $T^{+}_{i}$. Hence, there exist curves $\gamma_{a},\gamma_{b}\subset T^{+}_{i}$ from $s_{i}$ to $t_{i}$ on $P_{i}$ that together enclose $e_{i}$ and separate $e_{i}$ from $u$, see Figure 12.
Figure 12: An illustration for the proof of Claim 8: the curve $\gamma_{a}$ in $P_{i}$ separates $e_{i}$ from $u$.
Because $s_{i}$ and $t_{i}$ are visible from $u$, the curves intersect the segments $s_{i}u$ and $t_{i}u$ exactly in $s_{i}$ and $t_{i}$, respectively. We denote the curve contained in the triangle of $u,s_{i},t_{i}$ by $\gamma_{a}$. The curve $\gamma_{a}$ contains a point $x\in{o}$ because $s_{i}$ and $t_{i}$ lie in different components of $T^{+}_{i}-{o}$. Hence, the ray supporting $ux$ intersects $e_{i}$, i.e., $x>e_{i}$.
Next we show that if o{o} belong to $T_{i}$, then the curve $\gamma_{b}$ intersects o{o} in a point below $e_{i}$. Analogous to $\gamma_{a}$, the curve $\gamma_{b}$ also contains a point $x^{\prime}\in{o}$. It remains to show that $x^{\prime}$ is below $e_{i}$. Note that this does not hold, only in the case that $x^{\prime}$ is to the left of the ray $us_{i}$ or to the right of the ray $ut_{i}$; otherwise, $\gamma_{b}$ is below $e_{i}$. By Claim 7, the projection of the helper triangle separates $D$ from the vertices $s_{i},t_{i}$ in $D^{+}$. Consequently, the helper triangle $\triangle$ and $P_{i}$ intersect in two points $q_{1}$ and $q_{2}$. The points $q_{1}$ and $q_{2}$ must be above $e_{i}$, because $u$ forms a triangle with every edge of $\triangle$. By Claim 7 and the fact that the vertices of $o$ are contained in $D$, the vertices of $o$ and their convex hull are inside $\triangle$ in $D$, see Figure 13. Note that $o$ is not completely contained in $D$ because it contains an edge of the tunnel bottom.
Figure 13: An illustration for the proof of Claim 8. The wedge $W_{i}$ and the helper triangle $\triangle$ (yellow) intersect in the segment $q_{1}q_{2}$. We drew the projection of $D$ onto the the helper triangle. In $D$, the vertices of the sloop $o$ (purple) are contained in $\triangle$. Therefore, the point $x^{\prime}$ lies between the rays $uq_{1}$ and $uq_{2}$ on $W_{i}$ and thus below $e_{i}$.
This implies that $x^{\prime}$ is in a part of $\gamma_{b}$ that is bounded by the rays $uq_{1}$ and $uq_{2}$, see Figure 12. Therefore, $x^{\prime}$ lies below $e_{i}$ and is thus invisible from $u$. It follows that $x^{\prime}$ belongs to an invisible edge of o{o}, i.e., $x^{\prime}$ is contained in the edge of $o$ belonging to the tunnel bottom; we denote it by $b_{o}$. By Claim 6, no point of $b_{o}$ lies above $e_{i}$. Consequently, $x>e_{i}$ belongs to the tunnel roof, i.e., the rafter of o{o}.
Moreover, no triangle of $T_{i}$ incident to $b_{o}$ lies above $e_{i}$. Note that for each triangle in the bottom of $T_{i}$, there exists a tunnel loop o{o} such that $b_{o}$ is incident to it. Hence, no point of the bottom of $T_{i}$ lies above $e_{i}$, i.e., we obtain the following Property 1: the curve $\gamma_{a}$ does not contain points of the bottom of the tunnel $T_{i}$.
It remains to argue that there exists no further intersection point if ${o}\subset T_{i}$. We show that $(\gamma_{a}\cup\gamma_{b})\cap{o}=\{x,x^{\prime}\}$. Suppose that there exists a further point $x^{\prime\prime}\in(\gamma_{a}\cup\gamma_{b})\cap{o}$. Then o{o} must be contained in $P_{i}$ and we may assume that $x,x^{\prime\prime}$ are vertices of o{o}. However, then the edge $xx^{\prime\prime}$ intersects the closed curve formed by concatenating the segments $us_{i}$, $s_{i}t_{i}=e_{i}$, $t_{i}u$ – a contradiction to the properties of a geometric embedding. Consequently, $(\gamma_{a}\cup\gamma_{b})\cap{o}=\{x,x^{\prime}\}$. In particular, we obtain Property 2: $\gamma_{a}$ intersects each rafter of $T_{i}$ exactly once. This implies that $a_{i}$ intersects each rafter in $D$ exactly once and thus proves i).
Together, Properties 1 and 2 imply that $\gamma_{a}$ visits any section of $T_{i}$ exactly once. Consequently, the part of $a_{i}$ covered by the roof faces of any section forms an interval on $a_{i}$. This proves iii).
It remains to prove ii). By Property 2, $\gamma_{a}$ intersects each rafter exactly once. We denote the intersection points on the curve $\gamma_{a}$ with the rafters by $p^{\prime}_{1},\dots,p^{\prime}_{k}$ in the order they appear on $\gamma_{a}$, i.e., $p^{\prime}_{j}$ denotes the intersection of $\gamma_{a}$ with the $j$-th rafter of $T_{i}$. Let $p_{j}$ denote the projection of $p^{\prime}_{j}$ onto $S$, i.e., the intersection point of the $j$-th rafter with $a_{i}$ in $D\cup\mathcal{A}$. We show that $p_{1},\ldots,p_{k}$ appear in this order along $a_{i}$ in $D\cup\{a_{i}\}$, i.e., $p_{j}$ lies before $p_{j+1}$ for all $j$. To this end, we restrict our attention to the wedge $W_{i}\subset P_{i}$ bounded by the rays $uq_{1}$ and $uq_{2}$. For simplicity of the presentation, we transform $W_{i}$ such that (the interesting part of) $e_{i}$ is on the $x$-axis, $s_{i}$ left of $t_{i}$, and that the apex $u$ at infinity above, see also Figure 14(a).
(a) The curve $\gamma_{a}$ shields $e_{i}$ from $u$.
(b) $p^{\prime}_{j+1}$ lies below $\gamma^{\prime}$.
(c) $p^{\prime}_{j+1}$ lies above $\gamma^{\prime}$.
Figure 14: Illustration for the proof of Claim 8 ii).
Suppose for the purpose of a contradiction that $p_{j+1}$ lies left of $p_{j}$, see also Figures 14(b) and 14(c). Note that the initial subcurve of $\gamma_{a}$ until $p^{\prime}_{j}$, denoted by $\gamma_{a}^{\prime}$, shields the initial part of $e_{i}$ until $p_{j}$. If $p^{\prime}_{j+1}$ lies below $\gamma_{a}^{\prime}$ as depicted in Figure 14(b), then $p^{\prime}_{j+1}$ is not visible from $u$. However, by construction, all rafters and thus all points $p^{\prime}_{1},\dots,p^{\prime}_{k}$ are visible from the apex $u$. A contradiction. If $p^{\prime}_{j+1}$ lies above $\gamma_{a}^{\prime}$ as depicted in Figure 14(c), then $p^{\prime}_{i+1}$ is not visible from $u$ because of the subcurve $\gamma_{a}-\gamma_{a}^{\prime}$, a contradiction. Consequently, $p_{j}$ is left of $p_{j+1}$ for all $j$. ∎
In order to define a notion of left and right of $a_{i}$ in $D\cup\{a_{i}\}$, we enhance $D\cup\{a_{i}\}$ by the projection of $\triangle$. The arc $a_{i}$ partitions the interior of (the projection of) $\triangle$ into to two regions. Considering the orientation of $a_{i}$, it is easy to determine if a point within one of these two regions lies left or right.
Every drawing on the sphere can be transformed to a drawing in the plane where any face can be chosen as the outer face. In order to relate notions such as clockwise and counter-clockwise with our drawing on the sphere, we fix the face $f_{0}$ as the outer face of $D^{+}$ such that $v_{0},v_{1},v_{2}$ is a counter clockwise cycle. For $D$, these notions are then inherited because we consider a drawing $D$ that is an induced subdrawing of $D^{+}$.
Claim 9.
In $D\cup\{a_{i}\}$, the vertices of the left and right roof paths of $T_{i}$ are left and right of $a_{i}$.
Proof.
Let $u_{1},\dots,u_{k}$ denote the vertices of the left roof path, $v_{1},\dots,v_{k}$ the vertices of the central roof path, and $w_{1},\dots,w_{k}$ the vertices of the right roof path of $T_{i}$ in $D\cup\{a_{i}\}$. The value of $k$ depends on the number of tunnel sections of $T_{i}$, namely, k=2n4+ik=2n-4+i. For each $j=1,\ldots,k$, it holds that $u_{j},v_{j},w_{j}$ is a rafter which $a_{i}$ intersects exactly once by Claim 8 i). Consequently, if $u_{j}$ lies left of $a_{i}$ then $w_{j}$ lies right of $a_{i}$ and vice versa.
We consider a visible section of $T_{i}$. To this end, let $u_{j}u_{j+1}$ be a visible edge. We show that both $u_{j}$ and $u_{j+1}$ lie left of $a_{i}$ implying that $w_{j}$ and $w_{j+1}$ lie right of $a_{i}$. Because each vertex of the left or right roof path belongs to some visible section, this property implies the claim.
First consider the case that $u_{j}$ and $u_{j+1}$ lie right of $a_{i}$. By Claim 8 ii), the intersection $x_{j}$ of the rafter $u_{j},v_{j},w_{j}$ with $a_{i}$ lies before the intersection $x_{j+1}$ of the rafter $u_{j+1},v_{j+1},w_{j+1}$ with $a_{i}$. However, the cycle $u_{j},u_{j+1},v_{j+1},w_{j+1},w_{j},v_{j}$ is flipped, i.e., it has the outer face $f_{0}$ to its right, however it is supposed to lie to its left, see Figures 15(a) and 15(b). A contradiction.
Figure 15: Illustration for the proof of Claim 9.
Hence it remains to consider the case in which $u_{j}$ and $u_{j+1}$ lie on different sides. We consider the case that $u_{j}$ lies left and $u_{j+1}$ right of $a_{i}$ as illustrated in Figure 15(c). By crossing-freeness, exactly one of the edges $u_{j}u_{j+1}$ and $w_{j}w_{j+1}$ crosses $a_{i}$ between $x_{j}$ and $x_{j+1}$, while the other edge crosses before $x_{j}$ or after $x_{j+1}$ as depicted in Figure 15(c). However, then the roof faces of this section do not cover a consecutive part of $a_{i}$. A contradiction to Claim 8 iii). ∎
Claim 10.
In $\mathcal{A}$, for all $i,j$ with $i<j$, the arc $a_{j}$ crosses the arc $a_{i}$ from left to right.
Proof.
We restrict our attention to the vertices of the left and right roof paths of $T_{i}$ and $T_{j}$. Let $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\beta_{1},\beta_{2},\beta_{3}$ denote the vertices of two rafters of $T_{i}$ such that they are contained in the right and left roof path of $T_{j}$, respectively. For an illustration, consider Figure 16(a). We call their induced rafters, the $\alpha$-rafter and the $\beta$-rafter. By Claim 9, the vertices of a left/right roof path lie on the left/right of its corresponding arc.
(a) The vertices $\alpha_{1},\alpha_{2},\alpha_{3},\beta_{1},\beta_{2},\beta_{3}$.
(b) The arcs $a_{i}$ and $a_{j}$ cross.
(c) Arc $a_{j}$ cannot cross $a_{i}$ from right to left.
Figure 16: Illustration for the proof of Claim 10.
Suppose the arcs $a_{i}$ and $a_{j}$ do not cross. By symmetry, we may assume that $a_{j}$ lies left of $a_{i}$, see Figure 16(b). Recall that the vertices $\alpha_{3}$ and $\beta_{3}$ lie to the right side of $a_{i}$, but on different sides of $a_{j}$. A contradiction.
Now, suppose $a_{j}$ crosses $a_{i}$ from right to left as in Figure 16(c). Then each vertex among $\alpha_{1},\alpha_{3},\beta_{1},\beta_{3}$ lies in one of the four regions defined by $a_{i}$ and $a_{j}$. Moreover $a_{j}$ separates the $\alpha$-rafter of $T_{i}$ containing $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$ from the $\beta$-rafter containing $\beta_{1}$$\beta_{2}$, and $\beta_{3}$. This implies that $a_{i}$ intersects the $\beta$-rafter before the $\alpha$-rafter. A contradiction to Claim 8 ii). ∎
Claim 11.
In $\mathcal{A}$, the order of intersections on each arc $a_{i}$ is the same as for $\ell_{i}$ in $L^{*}$.
Proof.
We consider any three arcs $a_{i},a_{j},a_{k}$ with $i<j<k$. By Claim 10, the vertices $s_{i}$$s_{j}$$s_{k}$$t_{i}$$t_{j}$, and $t_{k}$ lie on the correct sides for all three arcs. It is therefore the case that the order of intersections is correct on one arc of a triple if and only if it is correct on all three arcs. Suppose for a contradiction that order of intersections on $a_{i}$ is not correct, i.e., the intersection of $a_{j}$ and $a_{k}$ lies on the wrong side of $a_{i}$. By symmetry, we consider the case that the intersection is left of $a_{i}$ while it is supposed to lie right of $a_{i}$, see Figure 17.
(a) The subgraph of $G$ induced by vertices of the three tunnels $T_{i},T_{j},T_{k}$.
(b) A wrong order of intersections implies that $a_{i}$ intersects the $\beta$-rafter before the $\alpha$-rafter.
Figure 17: Illustration for the proof of Claim 11.
Let $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\beta_{1},\beta_{2},\beta_{3}$ denote the vertices of two rafters of $T_{i}$ such that they are contained in the left roof path of $T_{j}$ and in the right roof path of $T_{k}$, respectively. For an illustration, consider Figure 17(a). Consequently, by Claim 9, $\alpha_{1},\alpha_{2},\alpha_{3}$ lie left of $a_{j}$ and $\beta_{1},\beta_{2},\beta_{3}$ lie right of $a_{k}$. This implies that the intersection of the $\beta$-rafter with $a_{i}$ lies before the intersection of the $\alpha$-rafter with $a_{i}$. A contradiction to Claim 8 ii). ∎
Claim 12.
In $\mathcal{A}$, all intersections lie in one hemisphere of $S$ bounded by $a_{0}$.
Proof.
@@
Consider the plane $P$ through the arc $a_{0}$ and apex $u$ and assume without loss of generality that $P$ is horizontal. By construction and Claim 11, the first crossing on each arc $a_{i}$ is with the arc $a_{0}$; all other crossings come afterwards. Moreover, Claim 10 ensures that each arc $a_{i}$ crosses $a_{0}$ from left to right. Consequently, all arcs start in one hemisphere of $S$ cut by $P$; let us suppose it is the top hemisphere. Then, the intersection of any two arcs $a_{i}$ and $a_{j}$ lie on the bottom hemisphere; otherwise the arcs contain both intersections of the supporting great circles. ∎
Let $P^{\prime}$ denote a plane obtained by shifting the plane $P$ through $a_{0}$ and $u$ to the bottom, see also Figure 11(c). Then, we ’project’ the arc arrangement onto the plane $P^{\prime}$ by considering the intersection of $P^{\prime}$ with the plane supporting the arc $a_{i}$. Clearly, this intersection yields a line. Hence, we obtain a line arrangement in the plane $P^{\prime}$ that inherits the combinatorics of $\mathcal{A}$ and thus, proves the stretchability of $L$. This finishes the proof of Theorem 4. ∎
@@
The constructed 2-complex in the proof of Theorem 4 was not pure. Specifically, the special edges are not contained in any triangle. We can obtain a pure 2-complex by adding one new vertex to each special edge such that it forms a special triangle. Similarly, we can add a private vertex to each triangle to form a pure 3-complex $C^{\prime}$. Given a geometric embedding of $C$, these new vertices can easily be added close enough to their defining set in $C$. Hence, $C$ has a geometric embedding if and only if $C^{\prime}$ has a geometric embedding in $\mathbb{R}^{3}$. Together with Theorem 4, this very small modification shows hardness of $\textsc{GEM}_{3\rightarrow 3}$.